OpenStudy (anonymous):

f(b)-f(a)/b-a, a=2, a=4, is their another way I'm supposed to be putting this in than just simply putting in the a and b values given: -2x^2-5x+4

OpenStudy (freckles):

I bet b is equal to one of those values

OpenStudy (freckles):

also to find f(4) or f(2) you need to use the function given

OpenStudy (anonymous):

So would the top line be -48-(-14) for f(b)-f(a)?

OpenStudy (anonymous):

Do you mean \(\frac{ f(b)-f(a) }{ b-a }\), where \( a = 2, b =4 \) and where \[f(x)=-2x^2-5x+4\] f(2) = -14 f(b) = -38 a - b = -2 so \[\frac{ f(4)-f(2) }{ 4-2 }= \frac{ -38 +14 }{ 2 } = -12\]

OpenStudy (anonymous):

but i was wrong you are right

OpenStudy (anonymous):

i don't think that is right

OpenStudy (freckles):

-48 is right

OpenStudy (freckles):

for f(4)

OpenStudy (anonymous):

yeah, my arithmetic is not good

OpenStudy (freckles):

i just wanted to say that because 2 people have already said it was something different :p

OpenStudy (anonymous):

so f(b) is -48 f(a) is -14

OpenStudy (anonymous):

yes sorry my arithmetic is louzy too

OpenStudy (anonymous):

especially since you need to be quick, otherwise someone else is jotting in anwers